# On the spectrum of closed neighborhood corona product of graph and its application

### v1

@@ Abstract
This paper introduces the concept of the closed neighborhood corona product of the graph. We explore the mathematical features of this product graph, specifically in terms of its spectral characteristics. We have calculated the characteristic polynomials of the adjacency, Laplacian, and signless Laplacian matrices. Moreover, we investigate the conditions under which two graphs are cospectral regarding this product. A significant portion of our study is dedicated to computing the Kirchhoff index, the number of spanning trees and the sequence of non-cospectral equienergetic product graphs. We also outline specific criteria that determine when the product graph is integral.
@@
Keywords: Closed neighborhood corona, Integral graph, Kirchhoff Index, Spanning trees, equienergetic graphs.
@@ 1 Introduction
In graph theory, a graph $\mathcal{G}$ is denoted as an ordered pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{i}:i=1,2,\ldots,n\}$ represents the set of vertices, and =$\mathcal{E}=\{e_{i}:i=1,2,\ldots,m\}$ represents the set of edges. The vertex degree $d(v_{i})$ is defined as the number of edges that are connected to the vertex $v_{i}$. Hence, the adjacency matrix $\mathcal{A}_{\mathcal{G}}=(a_{ij})_{n\times n}$ is a critical representation of the graph $\mathcal{G}$, where each entry $a_{ij}=1$ if there is an edge between vertex $v_{i}$ and vertex $v_{j}$ i.e., $v_{i}$ and $v_{j}$ are adjacent, and $a_{ij}=0$ otherwise. Additionally, two crucial matrices associated with a graph $\mathcal{G}$: the Laplacian matrix $\mathcal{L}_{\mathcal{G}}=\mathcal{D}_{\mathcal{G}}-\mathcal{A}_{\mathcal{G}}$ and the signless Laplacian matrix $\mathcal{Q}_{\mathcal{G}}=\mathcal{D}_{\mathcal{G}}+\mathcal{A}_{\mathcal{G}}$. The diagonal matrix $\mathcal{D}_{\mathcal{G}}$ has the $i^{th}$ diagonal entry as $d(v_{i})$, representing the degree of vertex $v_{i}$. The shortest distance between any two vertices $v_{i}$ and $v_{j}$ of the graph $\mathcal{G}$ is denoted as $d(v_{i},v_{j})$, and for any $v_{i}\in\mathcal{V}$, $N(v_{i})=\{v_{j}\in\mathcal{V}:v_{i}\text{ and }v_{j}\text{ are adjacent}\}$. The energy of a graph $\mathcal{G}$ of order $n$ is the sum of the absolute eigenvalues of the adjacency matrix, given by
@@
where $\lambda_{i}^{\prime}s$ are the eigenvalues of $\mathcal{A}_{\mathcal{G}}$. A graph $\mathcal{G}$ is classified as Integral if all the eigenvalues of its adjacency matrix are integers.. This discussion focuses exclusively on simple, undirected, and finite graphs.
The corona product for graphs was first defined by R Frucht and F Harary in 1970 [6]. Later, Cam McLeman and Erin McNicholas [11] introduced the coronal of graphs to obtain the adjacency spectra for arbitrary graphs. Shu and Gui [3] later generalized this concept and defined the coronal of the Laplacian and the signless Laplacian matrix of graphs. Since then, various researchers have explored the variants of corona product for both signed and unsigned graphs [2, 7, 8, 10, 13, 14, 15]. The present paper introduces the concept of the closed neighborhood corona product of graph. In this analysis, we compute the product graph’s adjacency, Laplacian, and signless Laplacian characteristic polynomials. We further establish a condition for the product graph to be cospectral. Additionally, we apply our findings to obtain the number of spanning trees, the Kirchhoff Index, sequence of non-cospectral equienergetic product graphs and the conditions required for the product graph to be Integral.
@@ 2 Preliminaries
Consider two matrices $\mathcal{A}=(a_{ij})$ with dimensions $m_{1}\times n_{1}$ and $\mathcal{B}=(b_{ij})$ with dimensions $m_{2}\times n_{2}$. The operation known as the Kronecker product, symbolized by $\mathcal{A}\otimes\mathcal{B}$, yields a new matrix with dimensions $m_{1}m_{2}\times n_{1}n_{2}$. This procedure involves taking each element $a_{ij}$ from matrix $\mathcal{A}$ and multiplying it by the entire matrix $\mathcal{B}$, effectively resizing $\mathcal{B}$ by the factor of $a_{ij}$. As highlighted by Neumaier (1992) [12], this technique allows for the assembly of a larger matrix from two smaller matrices.
@@ 2.1 Definition [11]
The coronal of Laplacian and signless Laplacian matrix was introduces by Shu and Gui [3] and is given by $\chi_{\mathcal{L}_{\mathcal{G}}}=1_{n}(xI_{n}-\mathcal{L}_{\mathcal{G}})^{-1}1_{n}$ and $\chi_{\mathcal{Q}_{\mathcal{G}}}=1_{n}(xI_{n}-\mathcal{Q}_{\mathcal{G}})^{-1}1_{n}$ respectively.
@@ 3 Spectral properties
(Closed neighborhood corona)  Consider two graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ with $n_{1}$ and $n_{2}$ vertices and $e_{1}$ and $e_{2}$ edges respectively. The closed neighborhood corona of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, denoted by $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$, is a new graph obtained by creating $n_{1}$ copies of $\mathcal{G}_{2}$. Each vertex of the $i^{th}$ copy of $\mathcal{G}_{2}$ is then connected to the $i^{th}$ vertex and neighborhood of the $i^{th}$ vertex ($u_{i}$) of $\mathcal{G}_{1}$.
@@
Figure 1: Closed neighborhood corona product of $C_{4}$ and $C_{3}$.
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{1}$ be arbitrary graphs on $n_{1}$ and $n_{2}$ vertices, respectively. Following the labelling given by I. Gopalapillai [7], we will label the vertices of $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$ as follows. Let $V(\mathcal{G}_{1})=\{u_{i}:i=1,2,\ldots,n_{1}\},$ Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be arbitrary graphs with $n_{1}$ and $n_{2}$ vertices, respectively. Following the vertex labeling approach introduced by I. Gopalapillai [7], we define the labeling of the vertices of the closed neighborhood corona product $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$. Denote the vertex sets of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ as $V(\mathcal{G}_{1})=\{u_{i}:i=1,2,\ldots,n_{1}\}$ and $V(\mathcal{G}_{2})=\{v_{i}:i=1,2,\ldots,n_{2}\}$, respectively. For each vertex $u_{i}$ in $\mathcal{G}_{1}$, we associate a copy of $\mathcal{G}_{2}$, denoted as the set $\{v_{1}^{i},v_{2}^{i},\ldots,v_{n_{2}}^{i}\}$, where $v_{j}^{i}$ represents the vertex in the $i^{th}$ copy of $\mathcal{G}_{2}$ corresponding to the vertex $v_{j}$ in $\mathcal{G}_{2}$. This labeling scheme facilitates the analysis of the product graph $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$ by systematically identifying the vertices across the two graph components. Denote
@@ 3.2 Laplacian spectrum
$\begin{split}\det[B_{L}]&=\det\Big{[}(x-n_{2}r_{1}-n_{2})I_{n_{1}}-\mathcal{L}_{\mathcal{G}_{1}}-\big{(}\textbf{1}^{T}_{n_{2}}\otimes(I_{n_{1}}+\mathcal{A}_{\mathcal{G}_{1}})\big{)}\{(x-r_{1}-1)I_{n_{2}}-\mathcal{L}_{\mathcal{G}_{2}}\}^{-1}\otimes I_{n_{1}}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \big{(}\textbf{1}_{n_{2}}\otimes(I_{n_{1}}+\mathcal{A}_{\mathcal{G}_{1}})\big{)}\Big{]}\\ &=\det\Big{[}(x-n_{2}r_{1}-n_{2})I_{n_{1}}-\mathcal{L}_{\mathcal{G}_{1}}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-r_{1}-1)(I_{n_{1}}+\mathcal{A}_{\mathcal{G}_{1}})^{2}\Big{]}\\ &=\prod_{i=1}^{n_{1}}\Big{[}x-n_{2}(r_{1}+1)-\gamma_{1i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-r_{1}-1)(1+\lambda_{1i})^{2}\Big{]}\\ &=\prod_{i=1}^{n_{1}}\Big{[}x-n_{2}(r_{1}+1)-\gamma_{1i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))(r_{1}+1-\gamma_{1i})^{2}\Big{]}\end{split}$
@@
$\begin{split}f(\mathcal{L}_{\mathcal{G}_{1}\boxtimes\mathcal{G}},x)&=\prod_{i=1}^{m}(x-(r_{1}+1)-\gamma_{i})^{n}\cdot\prod_{i=1}^{n}\Big{[}x-m(r_{1}+1)-\gamma_{1i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))(r_{1}+1-\gamma_{1i})^{2}\Big{]}\\ &=\prod_{i=1}^{m}(x-(r_{1}+1)-\gamma_{i})^{n}\cdot\prod_{i=1}^{n}\Big{[}x-m(r_{1}+1)-\gamma_{2i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))(r_{1}+1-\gamma_{2i})^{2}\Big{]}\\ &=f(\mathcal{L}_{\mathcal{G}_{2}\boxtimes\mathcal{G}},x).\end{split}$
@@
$\begin{split}f(\mathcal{L}_{\mathcal{G}\boxtimes\mathcal{G}_{1}},x)&=\prod_{i=1}^{n}(x-(r_{1}+1)-\gamma_{1i})^{m}\cdot\prod_{i=1}^{m}\Big{[}x-n(r_{1}+1)-\gamma_{i}-\chi_{\mathcal{L}_{\mathcal{G}_{1}}}(x-(r_{1}+1))(r_{1}+1-\gamma_{i})^{2}\Big{]}\\ &=\prod_{i=1}^{n}(x-(r_{1}+1)-\gamma_{2i})^{m}\cdot\prod_{i=1}^{m}\Big{[}x-n(r_{1}+1)-\gamma_{i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))(r_{1}+1-\lambda_{i})^{2}\Big{]}\\ &=f(\mathcal{L}_{\mathcal{G}\boxtimes\mathcal{G}_{2}},x).\end{split}$
@@ 3.3 Signless Laplacian spectrum
The later part of proof is same as Theorem 3.2. ∎
@@
Proof is same as Corollary 3.2.1. ∎
@@ 4.1 Kirchhoff Index
This renowned connection between the Laplacian spectrum and Kirchhoff index was determined in 1996 by Zhu et al. [16] and Gutman and Mohar [9].
@@
$\begin{split}&f(\mathcal{L}_{\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}},x)=\prod_{i=1}^{n_{2}}(x-(r_{1}+1)-\gamma_{2i})^{n_{1}}\cdot\prod_{i=1}^{n_{1}}\Big{[}x-n_{2}(r_{1}+1)-\gamma_{1i}-\frac{n_{2}}{x-r_{1}-1}(r_{1}+1-\gamma_{1i})^{2}\Big{]}\\ &=\prod_{i=1}^{n_{2}}(x-r_{1}-\gamma_{2i})^{n_{1}}\cdot\prod_{i=1}^{n_{1}}\frac{1}{x-r_{1}-1}\big{[}x^{2}-((r_{1}+1)(n_{2}+1)+\gamma_{1i})x+\gamma_{1i}(r_{1}+1)+n_{2}\gamma_{1i}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (2+2r_{1}-\gamma_{1i})\big{]}.\end{split}$ (5)

### v2

# On the spectrum of closed neighbourhood corona product of graph and its application

@@ Abstract
In this paper, we investigate the spectral properties of the closed neighborhood corona product of graphs, which was introduced by Harishchandra S. Ramane et al. [13] (cf. Polynomials Associated with Closed Neighborhood Corona and Neighborhood Complement Corona of Graphs). Based on their results, such as characteristic polynomials of the adjacency, Laplacian, and signless Laplacian matrices, we further investigate the spectral characteristics of this product graph. Specifically, we investigate conditions under which cospectrality occurs for this operation. Further, we determine the Kirchhoff index and count spanning trees and identify sequences of non-cospectral equienergetic product graphs. Finally, we develop criteria for when the product graph is integral and thereby contribute to a deeper understanding of the algebraic and combinatorial structure of the product graph.
@@
Keywords: Closed neighbourhood corona, Integral graph, Kirchhoff Index, Spanning trees, equienergetic graphs.
@@ 1 Introduction
In graph theory, a graph $\mathcal{G}$ is denoted as an ordered pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{i}:i=1,2,\ldots,n\}$ represents the set of vertices, and =$\mathcal{E}=\{e_{i}:i=1,2,\ldots,m\}$ represents the set of edges. The vertex degree $d(v_{i})$ is the number of edges connected to the vertex $v_{i}$. Hence, the adjacency matrix $\mathcal{A}_{\mathcal{G}}=(a_{ij})_{n\times n}$ is a critical representation of the graph $\mathcal{G}$, where each entry $a_{ij}=1$ if there is an edge between vertex $v_{i}$ and vertex $v_{j}$ i.e., $v_{i}$ and $v_{j}$ are adjacent, and $a_{ij}=0$ otherwise. Additionally, two crucial matrices associated with a graph $\mathcal{G}$: the Laplacian matrix $\mathcal{L}_{\mathcal{G}}=\mathcal{D}_{\mathcal{G}}-\mathcal{A}_{\mathcal{G}}$ and the signless Laplacian matrix $\mathcal{Q}_{\mathcal{G}}=\mathcal{D}_{\mathcal{G}}+\mathcal{A}_{\mathcal{G}}$. The diagonal matrix $\mathcal{D}_{\mathcal{G}}$ has the $i^{th}$ diagonal entry as $d(v_{i})$, representing the degree of vertex $v_{i}$. The shortest distance between any two vertices $v_{i}$ and $v_{j}$ of the graph $\mathcal{G}$ is denoted as $d(v_{i},v_{j})$, and for any $v_{i}\in\mathcal{V}$, $N(v_{i})=\{v_{j}\in\mathcal{V}:v_{i}\text{ and }v_{j}\text{ are adjacent}\}$. The energy of a graph $\mathcal{G}$ of order $n$ is the sum of the absolute eigenvalues of the adjacency matrix, given by
@@
where $\lambda_{i}^{\prime}s$ are the eigenvalues of $\mathcal{A}_{\mathcal{G}}$. A graph $\mathcal{G}$ is classified as Integral if all the eigenvalues of its adjacency matrix are integers. This discussion focuses exclusively on simple, undirected, and finite graphs.
Frucht and F Harary[6] first defined the corona product for graphs. Later, Cam McLeman and Erin McNicholas [11] introduced the coronal of graphs to obtain the adjacency spectra for arbitrary graphs. Shu and Gui [3] later generalized this concept and defined the coronal of the Laplacian and the signless Laplacian matrix of graphs. Since then, various researchers have explored the variants of corona product for both signed and unsigned graphs [2, 7, 8, 10, 14, 15, 16]. We establish a condition for the closed neighborhood corona product graph to be cospectral. Additionally, we apply our findings to obtain the number of spanning trees, the Kirchhoff Index, the sequence of non-cospectral equienergetic product graphs and the conditions required for the product graph to be Integral.
@@ 2 Preliminaries
Consider two matrices $\mathcal{A}=(a_{ij})$ with dimensions $m_{1}\times n_{1}$ and $\mathcal{B}=(b_{ij})$ with dimensions $m_{2}\times n_{2}$. The operation known as the Kronecker product, symbolized by $\mathcal{A}\otimes\mathcal{B}$, yields a new matrix with dimensions $m_{1}m_{2}\times n_{1}n_{2}$. This procedure involves taking each element $a_{ij}$ from matrix $\mathcal{A}$ and multiplying it by the entire matrix $\mathcal{B}$, effectively resizing $\mathcal{B}$ by the factor of $a_{ij}$. As highlighted by Neumaier (1992) [12], this technique allows for the assembling two smaller matrices.
@@ 2.1 Definition [11]
The coronal of Laplacian and signless Laplacian matrix was introduced by Gui [3] and is given by $\chi_{\mathcal{L}_{\mathcal{G}}}=1_{n}(xI_{n}-\mathcal{L}_{\mathcal{G}})^{-1}1_{n}$ and $\chi_{\mathcal{Q}_{\mathcal{G}}}=1_{n}(xI_{n}-\mathcal{Q}_{\mathcal{G}})^{-1}1_{n}$ respectively.
@@ 3 Spectral properties
(Closed neighbourhood corona)  Consider two graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ with $n_{1}$ and $n_{2}$ vertices and $e_{1}$ and $e_{2}$ edges respectively. The closed neighbourhood corona of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, denoted by $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$, is a new graph obtained by creating $n_{1}$ copies of $\mathcal{G}_{2}$. Each vertex of the $i^{th}$ copy of $\mathcal{G}_{2}$ is then connected to the $i^{th}$ vertex and neighbourhood of the $i^{th}$ vertex ($u_{i}$) of $\mathcal{G}_{1}$.
@@
[…]
Figure 1: Closed neighbourhood corona product of $C_{4}$ and $C_{3}$.
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{1}$ be arbitrary graphs on $n_{1}$ and $n_{2}$ vertices, respectively. Following the labelling given by I. Gopalapillai [7], we will label the vertices of $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$ as follows. Let $V(\mathcal{G}_{1})=\{u_{i}:i=1,2,\ldots,n_{1}\},$ Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be arbitrary graphs with $n_{1}$ and $n_{2}$ vertices, respectively. Following the vertex labelling approach introduced by I. Gopalapillai [7], we define the labelling of the vertices of the closed neighbourhood corona product $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$. Denote the vertex sets of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ as $V(\mathcal{G}_{1})=\{u_{i}:i=1,2,\ldots,n_{1}\}$ and $V(\mathcal{G}_{2})=\{v_{i}:i=1,2,\ldots,n_{2}\}$, respectively. For each vertex $u_{i}$ in $\mathcal{G}_{1}$, we associate a copy of $\mathcal{G}_{2}$, denoted as the set $\{v_{1}^{i},v_{2}^{i},\ldots,v_{n_{2}}^{i}\}$, where $v_{j}^{i}$ represents the vertex in the $i^{th}$ copy of $\mathcal{G}_{2}$ corresponding to the vertex $v_{j}$ in $\mathcal{G}_{2}$. The labelling scheme facilitates the analysis of the product graph $\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}$ by systematically identifying the vertices across the two graph components. Denote
@@ 3.1 Adjacency spectrum
Note: The above result has already been proved by Harishchandra S. Ramane et al. [13], we are adding this theorem for the sake of proper understanding of the further application.
@@ 3.2 Laplacian spectrum
$\begin{split}\det[B_{L}]&=\det\Big{[}(x-n_{2}r_{1}-n_{2})I_{n_{1}}-\mathcal{L}_{\mathcal{G}_{1}}-\big{(}\textbf{1}^{T}_{n_{2}}\otimes(I_{n_{1}}+\mathcal{A}_{\mathcal{G}_{1}})\big{)}\{(x-r_{1}-1)I_{n_{2}}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\mathcal{L}_{\mathcal{G}_{2}}\}^{-1}\otimes I_{n_{1}}\big{(}\textbf{1}_{n_{2}}\otimes(I_{n_{1}}+\mathcal{A}_{\mathcal{G}_{1}})\big{)}\Big{]}\\ &=\det\Big{[}(x-n_{2}r_{1}-n_{2})I_{n_{1}}-\mathcal{L}_{\mathcal{G}_{1}}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-r_{1}-1)(I_{n_{1}}+\mathcal{A}_{\mathcal{G}_{1}})^{2}\Big{]}\\ &=\prod_{i=1}^{n_{1}}\Big{[}x-n_{2}(r_{1}+1)-\gamma_{1i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-r_{1}-1)(1+\lambda_{1i})^{2}\Big{]}\\ &=\prod_{i=1}^{n_{1}}\Big{[}x-n_{2}(r_{1}+1)-\gamma_{1i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))(r_{1}+1-\gamma_{1i})^{2}\Big{]}\end{split}$
@@
Note: The above result has already been proved by Harishchandra S. Ramane et al. [13], we are adding this theorem for the sake of proper understanding of the further application.
@@
$\begin{split}f(\mathcal{L}_{\mathcal{G}_{1}\boxtimes\mathcal{G}},x)&=\prod_{i=1}^{m}(x-(r_{1}+1)-\gamma_{i})^{n}\cdot\prod_{i=1}^{n}\Big{[}x-m(r_{1}+1)-\gamma_{1i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (r_{1}+1-\gamma_{1i})^{2}\Big{]}\\ &=\prod_{i=1}^{m}(x-(r_{1}+1)-\gamma_{i})^{n}\cdot\prod_{i=1}^{n}\Big{[}x-m(r_{1}+1)-\gamma_{2i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (r_{1}+1-\gamma_{2i})^{2}\Big{]}\\ &=f(\mathcal{L}_{\mathcal{G}_{2}\boxtimes\mathcal{G}},x).\end{split}$
@@
$\begin{split}f(\mathcal{L}_{\mathcal{G}\boxtimes\mathcal{G}_{1}},x)&=\prod_{i=1}^{n}(x-(r_{1}+1)-\gamma_{1i})^{m}\cdot\prod_{i=1}^{m}\Big{[}x-n(r_{1}+1)-\gamma_{i}-\chi_{\mathcal{L}_{\mathcal{G}_{1}}}(x-(r_{1}+1))\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (r_{1}+1-\gamma_{i})^{2}\Big{]}\\ &=\prod_{i=1}^{n}(x-(r_{1}+1)-\gamma_{2i})^{m}\cdot\prod_{i=1}^{m}\Big{[}x-n(r_{1}+1)-\gamma_{i}-\chi_{\mathcal{L}_{\mathcal{G}_{2}}}(x-(r_{1}+1))\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (r_{1}+1-\lambda_{i})^{2}\Big{]}\\ &=f(\mathcal{L}_{\mathcal{G}\boxtimes\mathcal{G}_{2}},x).\end{split}$
@@ 3.3 Signless Laplacian spectrum
The later part of the proof is the same as Theorem 3.2. ∎
@@
The proof is the same as Corollary 3.2.1. ∎
@@ 4.1 Kirchhoff Index
This renowned connection between the Laplaciathe n spectrum and Kirchhoff index was determined in 1996 by Zhu et al. [17] and Gutman and Mohar [9].
@@
$\begin{split}&f(\mathcal{L}_{\mathcal{G}_{1}\boxtimes\mathcal{G}_{2}},x)=\prod_{i=1}^{n_{2}}(x-(r_{1}+1)-\gamma_{2i})^{n_{1}}\cdot\prod_{i=1}^{n_{1}}\Big{[}x-n_{2}(r_{1}+1)-\gamma_{1i}-\frac{n_{2}}{x-r_{1}-1}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (r_{1}+1-\gamma_{1i})^{2}\Big{]}\\ &=\prod_{i=1}^{n_{2}}(x-r_{1}-\gamma_{2i})^{n_{1}}\cdot\prod_{i=1}^{n_{1}}\frac{1}{x-r_{1}-1}\big{[}x^{2}-((r_{1}+1)(n_{2}+1)+\gamma_{1i})x+\gamma_{1i}(r_{1}+1)\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +n_{2}\gamma_{1i}(2+2r_{1}-\gamma_{1i})\big{]}.\end{split}$ (5)